koenko:

y+2=5 how do i solve this

supie:

So basically, you have to find what number plus two equals 5.

supie:

\(?+2=5\) Which is the same thing as \(5-2=?\) Or \(y+2−2=5−2\) (subtracting `2` from both sides)

koenko:

thank u that helped

supie:

np

supie:

lmk when you get your answer and what you get as your answer

koenko:

ok